Aspire Faculty ID #16182 · Topic: NIMCET 2011 · Just now
NIMCET 2011

An anti-aircraft gun fires at a plane. Probabilities of hitting at slots 1,2,3,4 are $0.4,;0.3,;0.2,;0.1$. Probability that the gun hits the plane is

Solution

Probability of miss in all four shots: $(1-0.4)(1-0.3)(1-0.2)(1-0.1)$ $=0.6 \times 0.7 \times 0.8 \times 0.9$ $=0.3024$ Probability of at least one hit: $1-0.3024 = 0.6976$

Previous 10 Questions — NIMCET 2011

Nearest first

Next 10 Questions — NIMCET 2011

Ascending by ID
Ask Your Question or Put Your Review.

loading...